OpenStudy (anonymous):

how to solve (x+4)(x-15)=6x

OpenStudy (anonymous):

\[\left( x+4 \right)\left( x-15 \right)=6x\] \[x ^{2}-15x+4x-60=6x\] \[x ^{2}-11x-60-6x=0\] \[x ^{2}-17x-60=0\] \[x ^{2}-20x+3x-60=0\] factorise and find the solutions.